OpenStudy (snowcrystal):

Can someone help me please!!! I will fan and medal!! (Factoring: question below)

OpenStudy (snowcrystal):

\[x^2 +14x-72=0\]

OpenStudy (ipwnbunnies):

Two factors that multiply to give you -72 and add to give you 14 is -4 and 18. Can you go from there?

OpenStudy (snowcrystal):

umm i think so

OpenStudy (snowcrystal):

thank u

OpenStudy (snowcrystal):

wait how did u get -4 and 18?

OpenStudy (ipwnbunnies):

-4 and 18 are two factors of 72. -4*18 = -72 && -4+18 = 14. So I found the numbers I need when I expand the equation.

OpenStudy (snowcrystal):

okay. what do i do about the x?

OpenStudy (ipwnbunnies):

Well that's why you're factoring it out. You'll have (x + a)(x +b) = 0. With a question like this, I would assume you know how to factor quadratic equations. I can't explain factoring that well.

OpenStudy (ipwnbunnies):

So, when you factor your original equation, you'll get: (x - 4)(x + 18) = 0. You can FOIL it to check if you got your original equation.

OpenStudy (ipwnbunnies):

a & b are the two numbers I found that will add to -14 and multiply to -72.

OpenStudy (snowcrystal):

okay thanks. I am not great with any kind of factoring. its been a struggle to learn it.

OpenStudy (ipwnbunnies):

It's ok. If you just follow that process for your problems, you can get it. Find two numbers that add up to make the middle term and multiply to get the last term.

OpenStudy (snowcrystal):

thanks